### v2

# LANCE: Low Rank Activation Compression for Efficient On-Device Continual Learning

Abstract
[…]
1 Introduction
[…]
Several approaches have been proposed to address this issue. Checkpointing [chen2016checkpoint] reduces memory usage at the cost of extra recomputation; reversible networks [gomez2017reversible] eliminate the need to store activations but require specialized architectures; and quantization or activation pruning [chen2021actnn, barley2023activation, yu2022backrazor] compress or sparsify activations but may introduce approximation errors or hardware overhead. Alternatively, bio-inspired methods have explored efficient replacements for BP, such as forward-forward methods [dellaferrera2022pepita, hinton2022forward] and local learning rules [lillicrap2014random, nokland2016DFA, apolinario2025lls, Frenkel2021LearningNetworks, apolinario2025tess]. However, they usually incur accuracy degradation. More recently, low-rank approaches exploit redundancy in activations using SVD or Tucker decompositions [nguyen2024activation, nguyen2025beyond], but these typically recompute decompositions at each training step, adding significant overhead and making them unsuitable for continual learning, where repeated factorization across tasks would require storing large task-specific matrices.
A key insight from orthogonal gradient projection methods in continual learning [Saha2021GradientLearning, liang2024inflora, apolinario2024code] is that neural activations often lie in a stable, low-rank subspace that can be reused across training. This suggests a natural question: Can we identify a compact activation subspace and reuse it throughout fine-tuning, while simultaneously enabling efficient continual learning?
With the above in mind, we introduce LANCE (Low-rank Activation Compression), a framework that applies one-time higher-order singular value decomposition (HOSVD) to extract a reusable low-rank activation subspace, as shown in Figure˜1b. Unlike prior iterative approaches [nguyen2024activation, nguyen2025beyond], LANCE computes the decomposition only once at the beginning of training. Subsequent activations are projected onto this fixed subspace throughout training, avoiding repeated decompositions, thereby significantly reducing memory usage, computational cost, and hardware overhead. In contrast to system-level methods such as TinyTL [lin2020tinytl] or the 256KB training engine [lin2022ondevice], which freeze layers or restrict updates to fit memory budgets, LANCE enables full backpropagation across all layers by directly compressing stored activations. Importantly, these fixed subspaces also allow assigning new tasks to orthogonal components, inherently supporting on-device continual learning without storing large task-specific matrices as in previous continual learning (CL) methods [Saha2021GradientLearning, Saha2023ContinualProjection, apolinario2024code, Lin2022TRGP:Learning, Lin22CUBER, SD2023, lmsp2025]. By eliminating repeated decompositions while preserving full-model fine-tuning, LANCE provides a practical and scalable solution for memory- and energy-efficient on-device continual learning.
Our main contributions are summarized as follows:
  • We propose LANCE, a one-shot activation compression framework that constructs a reusable low-rank subspace via HOSVD, avoiding repeated decompositions.
  • We show that LANCE enables both efficient fine-tuning and continual learning on edge devices by projecting activations onto fixed low-rank subspaces, which can be partitioned across tasks to mitigate forgetting.
  • We empirically validate LANCE on fine-tuning benchmarks (CIFAR-10/100, Pets, Flowers102, CUB-200 with MCUNet, MobileNetV2, ResNet18/34) and continual learning benchmarks (Split CIFAR-100, Split MiniImageNet, 5-Datasets), demonstrating memory savings of up to $250\times$ while maintaining accuracy within $2\%$ of full backpropagation and achieving competitive performance with full-rank gradient projection methods at a fraction of the memory cost.
2 Methodology
In this section, we outline our methodology. We first describe the main memory bottleneck for on-device learning with full backpropagation (BP), then review higher-order singular value decomposition (HOSVD), and finally present our proposed method, including an overview, step-by-step pseudocode, a complexity analysis, and a convergence analysis.
[…]
2.1 Preliminaries
Training deep neural networks requires storing intermediate activations during the forward pass so that they can be reused for gradient computation in backpropagation, as illustrated in Figure˜1a. Let $f({\mathbf{x}};{\bm{\theta}})$ denote a neural network with input $\mathbf{x}$} and parameters $\bm{\theta}$}. During forward propagation, each layer $l$ produces an activation tensor ${\bm{\mathsfit{X}}}^{(l)}$ that is stored in SRAM for later use in computing the gradient: $\nabla_{{\bm{\theta}}^{(l)}}\mathcal{L}\;=\;\frac{\partial\mathcal{L}}{\partial{\bm{\mathsfit{X}}}^{(l)}}\;\frac{\partial{\bm{\mathsfit{X}}}^{(l)}}{\partial{\bm{\theta}}^{(l)}}$ , where $\mathcal{L}$ is the loss function. The cumulative memory footprint of storing $\{{\bm{\mathsfit{X}}}^{(l)}\}_{l=1}^{L}$ across all $L$ layers often dominates SRAM usage, especially in convolutional networks. For example, on CIFAR-100 with ResNet18, storing activations during training can consume up to $5\times$ more memory than storing weights, quickly exhausting the limited resources of edge devices. A natural way to mitigate this bottleneck is to exploit the low-rank structure of activations. Instead of storing the full tensor ${\bm{\mathsfit{X}}}^{(l)}$, we can approximate it with a compressed representation and reconstruct (or use) it during backpropagation. While classical singular value decomposition (SVD) is defined for matrices, activations are inherently higher-order (e.g., batch $\times$ channels $\times$ height $\times$ width), making tensor decompositions more suitable.
Higher-Order Singular Value Decomposition (HOSVD).
HOSVD (or $N$-mode SVD) provides a principled way to decompose a tensor into mode-specific subspaces. Given an order-$d$ activation tensor ${\bm{\mathsfit{X}}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$, HOSVD expresses it as ${\bm{\mathsfit{X}}}\;=\;{\bm{\mathsfit{G}}}\times({\bm{U}}_{1},{\bm{U}}_{2},\cdots,{\bm{U}}_{d}),$ where ${\bm{\mathsfit{G}}}\in{\mathbb{R}}^{r_{1}\times r_{2}\times\cdots\times r_{d}}$ is the core tensor, ${\bm{U}}_{i}\in{\mathbb{R}}^{n_{i}\times r_{i}}$ are orthonormal factor matrices for each mode $i$, and $\times$ denotes the tensor–matrix product. The ranks $\{r_{i}\}_{i=1}^{d}$ control the compression level, with $r_{i}\ll n_{i}$ in the low-rank regime. To compute HOSVD, the tensor $\bm{\mathsfit{X}}$} is unfolded along each mode $i$ into a matrix ${\bm{X}}_{i}\in{\mathbb{R}}^{n_{i}\times\prod_{j\neq i}n_{j}}$, followed by a matrix SVD ${\bm{X}}_{i}={\bm{U}}_{i}{\bm{\Sigma}}_{i}{\bm{V}}_{i}^{\top}$. Truncating to the top-$r_{i}$ singular vectors yields ${\bm{U}}^{(i)}$, which captures the dominant subspace of mode $i$. Thus, HOSVD produces a compact multi-linear approximation of $\bm{\mathsfit{X}}$} that preserves most of its structure while discarding redundancy.
2.2 Proposed Method: LANCE
Algorithm 1 LANCE: Low-Rank Activation Compression for On-device Continual Learning
0: Pretrained network $f({\mathbf{x}};{\bm{\theta}})$ with layers $l{=}1,\dots,L$; calibration batches $\{\mathcal{B}_{j}\}_{j=1}^{N}$; energy thresholds $\varepsilon$ (LANCE) and $\varepsilon_{\text{CL}}$ (CL); (CL only) memory ${\bm{M}}^{t-1}$ from previous task
1:Notation: we omit layer superscripts for readability; all steps are applied per layer.
2:Phase I: One-Shot Subspace Calibration (offline)
3:for each mode $i\in\{1,\dots,d\}$ do
4:  Estimate covariance ${\bm{B}}_{i}$ from calibration activations via Equation˜2
5:  if i=di{=}d then
6:   $\hat{{\bm{B}}}_{d}\leftarrow({\bm{I}}-{\bm{M}}^{t-1}{\bm{M}}^{t-1\top})\,{\bm{B}}_{d}$ // For non-CL or $t=1$, ${\bm{M}}^{t-1}$ is a zero matrix
7:   Compute SVD, $\hat{{\bm{B}}}_{d}=\hat{{\bm{U}}}_{d}\hat{{\bm{\Sigma}}}_{d}\hat{{\bm{U}}}_{d}^{\top}$, and ${\bm{B}}_{d}={\bm{U}}_{d}{\bm{\Sigma}}_{d}{\bm{U}}_{d}^{\top}$
8:   Retain top $r_{d}$ columns of $\hat{{\bm{U}}}_{d}$ satisfying Equation˜5; set ${\bm{U}}_{d}\leftarrow\hat{{\bm{U}}}_{d}[:,1\!:\!r_{d}]$
9:  else
10:   Compute SVD ${\bm{B}}_{i}={\bm{U}}_{i}{\bm{\Sigma}}_{i}{\bm{U}}_{i}^{\top}$
11:   Retain top $r_{i}$ columns of ${\bm{U}}_{i}$ satisfying Equation˜3
12:  end if
13:end for
14:Phase II: Fine-Tuning with Low-Rank Activations (on-device)
15:for each minibatch $\mathcal{B}$ do
16:  Forward: project ${\bm{X}}^{(l)}$ to cores ${\bm{\mathsfit{G}}}^{(l)}={\bm{\mathsfit{X}}}^{(l)}\times({\bm{U}}_{1}^{\top},\dots,{\bm{U}}_{d}^{\top})$; store ${\bm{\mathsfit{G}}}^{(l)}$ only
17:  Backward: compute $\nabla_{{\bm{\theta}}^{(l)}}\mathcal{L}$ using ${\bm{\mathsfit{G}}}^{(l)}\times({\bm{U}}_{1},\dots,{\bm{U}}_{d})$; update ${\bm{\theta}}^{(l)}$
18:end for
19:Phase III (CL only): Memory Construction & Update (offline, after task $t$)
20: Estimate last-mode covariance ${\bm{B}}^{t}_{d}$ from $N_{\text{CL}}$ calibration mini-batches as in Equation˜2
21: Compute $\hat{{\bm{B}}}^{t}_{d}\leftarrow({\bm{I}}-{\bm{M}}^{t-1}{\bm{M}}^{t-1\top})\,{\bm{B}}^{t}_{d}$, then SVD on $\hat{{\bm{B}}}^{t}_{d}$ and ${\bm{B}}^{t}_{d}$
[…]
To address the challenges of training DNNs on low-power devices, we propose LANCE, a one-shot HOSVD-based low-rank activation compression method for efficient on-device learning. LANCE directly targets the primary memory bottleneck in backpropagation: the need to store intermediate activations ${\bm{\mathsfit{X}}}^{(l)}$ from hidden layers. By projecting activations into a compact low-rank subspace, LANCE reduces both the memory required to store ${\bm{\mathsfit{X}}}^{(l)}$ and the number of computations needed to backpropagate through them.
Formally, given projection matrices $\{{\bm{U}}_{i}\}_{i=1}^{d}$ obtained via HOSVD, each activation tensor ${\bm{\mathsfit{X}}}^{(l)}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ is compressed into a smaller core tensor ${\bm{\mathsfit{G}}}^{(l)}\in{\mathbb{R}}^{r_{1}\times r_{2}\times\cdots\times r_{d}}$:
${\bm{\mathsfit{G}}}^{(l)}\;=\;{\bm{\mathsfit{X}}}^{(l)}\times({\bm{U}}_{1}^{(l)\top},{\bm{U}}_{2}^{(l)\top},\cdots,{\bm{U}}_{d}^{(l)\top}).$ (1)
The compressed ${\bm{\mathsfit{G}}}^{(l)}$ is saved instead of the full activation and reused during backpropagation. Unlike prior works that recompute low-rank decompositions every iteration, LANCE computes the projection matrices only once at the beginning of fine-tuning and reuses them across all epochs. This one-shot approach avoids expensive per-iteration SVDs and is particularly well-suited for energy-constrained devices. The method, summarized in Algorithm 1, has two phases: (1) a one-shot subspace calibration performed offline, and (2) on-device fine-tuning with low-rank activations, which projects activations into low-rank subspaces during the forward pass and computes gradients during the backward pass using the compressed activations (Figure˜1b). Each step is detailed below.
Phase I: One-Shot Subspace Calibration (offline).
We estimate the subspace for each hidden layer $l$ using activations from $N$ calibration mini-batches passed through the pretrained model. For each mode $i$, we maintain a running covariance estimate:
${\bm{B}}_{i}[t]\;=\;\tfrac{1}{t}\Big((t-1){\bm{B}}_{i}[t-1]\;+\;\tfrac{1}{\prod_{j\neq i}n_{j}}{\bm{X}}_{i}^{(l)}{\bm{X}}_{i}^{(l)\top}\Big),$ (2)
with ${\bm{B}}_{i}[0]=0$, where ${\bm{X}}_{i}^{(l)}$ is the mode-$i$ unfolding of ${\bm{\mathsfit{X}}}^{(l)}$. After $N$ batches, we compute an SVD decomposition, ${\bm{B}}_{i}={\bm{U}}_{i}{\bm{\Sigma}}_{i}{\bm{U}}_{i}^{\top}$. We then truncate to the top-$r_{i}$ singular vectors according to an energy threshold $\varepsilon$:
$\frac{\sum_{j=1}^{r_{i}}\sigma_{j}}{\sum_{j=1}^{n_{i}}\sigma_{j}}\;\geq\;\varepsilon,$ (3)
where $\sigma_{j}$ is the $j$-th singular value of ${\bm{B}}_{i}$. This ensures that the subspace captures at least an $\varepsilon$ fraction of activation variance. Because this decomposition is computed only once (before fine-tuning) and amortized across training, its cost is negligible relative to repeated decompositions.
Phase II: Fine-Tuning with Low-Rank Activations (on-device).
During fine-tuning with BP (Figure˜1b), we distinguish the forward and backward passes.
Forward pass. Each activation ${\bm{\mathsfit{X}}}^{(l)}$ is projected into the learned low-rank subspace using the precomputed matrices $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$, following Equation˜1 to obtain a core tensor ${\bm{\mathsfit{G}}}^{(l)}$. Only the compact tensor ${\bm{\mathsfit{G}}}^{(l)}$ and the fixed $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$ need to be stored for backpropagation. This reduces the memory footprint from $\mathcal{O}(\prod_{i=1}^{d}n_{i})$ to $\mathcal{O}(\prod_{i=1}^{d}r_{i}+\sum_{i=1}^{d}n_{i}r_{i})$ per activation.
Backward pass. The gradient with respect to the weights is computed using the compressed activations. Given the gradient of the loss with respect to the layer output, $\nabla_{{\bm{\mathsfit{X}}}^{(l+1)}}\mathcal{L}$, we reconstruct the contribution with respect to the layer’s parameters ${\bm{\theta}}^{(l)}$ by reversing the projections:
$\nabla_{{\bm{\theta}}^{(l)}}\tilde{\mathcal{L}}\;=\;\frac{\partial\mathcal{L}^{\top}}{\partial{\bm{\mathsfit{X}}}^{(l+1)}}\;\big({\bm{\mathsfit{G}}}^{(l)}\times({\bm{U}}^{(l)}_{1},\cdots,{\bm{U}}^{(l)}_{d})\big)$. (4)
Because $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d}$ are orthonormal, this operation preserves gradient directions up to the truncation error introduced by the low-rank approximation. We prove that the approximated gradient $\nabla_{{\bm{\theta}}^{(l)}}\tilde{\mathcal{L}}$ is a descent direction and that the overall method converges in Sec. 2.4. Finally, gradients with respect to activations, $\nabla_{{\bm{\mathsfit{X}}}^{(l)}}\mathcal{L}$, are computed as usual by automatic differentiation; the compression only affects which forward tensors are stored (cores ${\bm{\mathsfit{G}}}^{(l)}$ vs. full ${\bm{\mathsfit{X}}}^{(l)}$), not the functional form of the backward.
2.3 LANCE for On-Device Continual Learning
We extend LANCE to the continual learning (CL) setting, where tasks arrive sequentially $t=1,2,\dots,T$. After completing task $t$, we retain for each layer a compact set of important directions on the last mode (the input dimension of the layer’s weights), denoted by ${\bm{M}}^{t}$. When calibrating the HOSVD for the next task t+1t{+}1, the new columns of ${\bm{U}}_{d}$ (the mode-$d$ factors) are chosen to lie strictly in the null space of ${\bm{M}}^{t}$. This ensures that task-t+1t{+}1 learns from directions orthogonal to those already used by previous tasks, thereby preventing interference and mitigating forgetting.
For clarity, we omit the explicit layer superscript in the following notation; all operations (memory construction, calibration, and updates) are applied independently at each layer.
Phase III: Memory construction and update.
At the end of training on task $t$, we estimate the subspace ${\bm{M}}^{t}$ by accumulating the covariance of activations along the last mode, ${\bm{B}}_{d}^{t}$, as in Equation˜2 using $N_{\text{CL}}$ calibration mini-batches. An SVD decomposition of ${\bm{B}}^{t}_{d}$ yields a basis of principal directions, and we retain the top $r_{d}$ components that capture at least an energy fraction $\varepsilon_{\text{CL}}$. To ensure orthogonality with previously stored directions, we project ${\bm{B}}^{t}_{d}$ onto the complement of ${\bm{M}}^{t-1}$: $\hat{\bm{B}}^{t}_{d}\;=\;{\bm{B}}^{t}_{d}-{\bm{M}}^{t-1}{\bm{M}}^{t-1\top}{\bm{B}}^{t}_{d}$, where ${\bm{M}}^{0}$ is initialized as empty. We select the smallest $r_{d}$ such that:
$\frac{\sum_{j=1}^{n_{d}}\sigma_{j}-\sum_{j=r_{d}}^{n_{d}}\hat{\sigma}_{j}}{\sum_{j=1}^{n_{d}}\sigma_{j}}\;\geq\;\varepsilon_{\text{CL}}.$ (5)
[…]
Null-space constrained calibration.
For a new task t+1t{+}1, we update the one-shot calibration of LANCE by incorporating the memory constraint only in mode-$d$; all other modes follow the procedure in Section 2.2. We compute the covariance ${\bm{B}}_{d}$ for the new task, remove contributions along ${\bm{M}}^{t}$, $\hat{\bm{B}}_{d}\;=\;{\bm{B}}_{d}-{\bm{B}}_{d}{\bm{M}}^{t}{\bm{M}}^{t\top}$, and select the top $r_{d}$ eigenvectors of $\hat{\bm{B}}_{d}$ that satisfy an energy threshold $\varepsilon$ similar to Equation˜5. The resulting low-rank factors used during fine-tuning are $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{d-1}\cup\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$, where $\hat{\bm{U}}^{(l)}_{d,[:,1:r_{d}]}$ are the retained singular vectors of $\hat{\bm{B}}_{d}$. This guarantees that activations of the new task are compressed into a subspace orthogonal to those of previous tasks, thereby reducing catastrophic forgetting.
Discussion.
Both memory update and null-space constrained calibration are performed offline, so they do not affect the efficiency of on-device fine-tuning. The method requires storing only the compact memory matrices $\{{\bm{M}}^{(l)}\}$ and applying inexpensive projections during calibration. By construction, interference with past tasks is eliminated along the stored subspaces, and forgetting is mitigated without replay buffers or large task-specific models. In contrast to prior gradient-projection methods [Saha2021GradientLearning, Saha2023ContinualProjection, Lin2022TRGP:Learning, Lin22CUBER, apolinario2024code], which explicitly use $\{{\bm{M}}^{(l)}\}$ for online gradient projection and must store such matrices in SRAM during training, our approach achieves continual learning naturally by projecting activations into fixed low-rank subspaces that lie in the null space of previous tasks, retaining LANCE’s memory efficiency.
2.4 Convergence Analysis
We provide a brief convergence analysis showing that LANCE yields valid descent directions and converges to projected stationary points. Full proofs are deferred to Appendix B.
Assumption A1 (Orthogonal truncation)
For every layer $l$ and mode $i$, the calibration bases ${\bm{U}}_{i}^{(l)}$ have orthonormal columns; hence $\mathsf{P}^{(l)}:={\bm{U}}_{1}^{(l)}{\bm{U}}_{1}^{(l)\top}\otimes\cdots\otimes{\bm{U}}_{d}^{(l)}{\bm{U}}_{d}^{(l)\top}$ is an orthogonal projector.
Assumption A2 (Smoothness)
The empirical loss $\mathcal{L}({\bm{\theta}})$ is $L$-smooth.
Assumption A3 (Layer linearity in inputs)
Each trainable layer is linear in its input during backpropagation (dense or convolutional after unfolding). Nonlinearities are piecewise linear and treated as fixed locally.
Assumption A4 (Energy capture)
One-shot HOSVD retains a fraction $\varepsilon\in(0,1]$ of activation energy per mode, inducing a bound on gradient leakage outside the retained input subspace (used in Proposition˜1).
Theorem 2.1(Projected gradient & descent)
Under Assumptions A1 and A3, for any layer $l$ with input projector $\mathsf{P}^{(l)}$, the LANCE weight gradient equals the right-Frobenius orthogonal projection of the full gradient: $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}=\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\,\mathsf{P}^{(l)}.$ Consequently, $\langle\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}},\,\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\rangle=\|\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}\|_{F}^{2}\geq 0$, so $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it vanishes.
Theorem 2.2(Monotone decrease & projected stationarity)
Let $\mathsf{Q}$ be the block projector that right-multiplies each layer’s weight gradient by its $\mathsf{P}^{(l)}$. Under Assumption A2, the LANCE update ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta\,\mathsf{Q}\,\nabla\mathcal{L}({\bm{\theta}}_{k})$ with $\eta\in(0,1/L]$ satisfies $\mathcal{L}({\bm{\theta}}_{k+1})\;\leq\;\mathcal{L}({\bm{\theta}}_{k})\;-\;\tfrac{\eta}{2}\,\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2},$ hence $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$ and every limit point $\bm{\theta}$$_{\star}$ obeys $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$.
Proposition 1(Stationarity gap vs. truncation)
If along the iterates $\|(I-\mathsf{Q})\nabla\mathcal{L}({\bm{\theta}})\|_{2}\leq C\sqrt{1-\varepsilon}$ for some $C>0$, then any limit point $\bm{\theta}$$_{\star}$ satisfies $\|\nabla\mathcal{L}({\bm{\theta}}_{\star})\|_{2}\leq C\sqrt{1-\varepsilon}$.
Insight.
Together, these results show that LANCE behaves like full BP restricted to the low-rank activation subspace: updates are valid descent directions, the loss decreases monotonically, and iterates converge to projected stationary points. The residual gap depends only on the energy threshold $\varepsilon$, so higher $\varepsilon$ yields solutions closer to true stationary points, while lower $\varepsilon$ trades accuracy for efficiency in a quantifiable way.
3 Experimental Evaluation
We evaluate LANCE in two complementary settings: (i) single-task fine-tuning, where a pretrained model is adapted to a target dataset on an edge device, and (ii) continual learning, where multiple tasks arrive sequentially and must be learned without catastrophic forgetting. Our experiments aim to answer the following questions:
  • Q1
    Does the one-shot subspace calibration preserve gradient fidelity in practice?
  • Q2
    How effective is LANCE at reducing activation memory and computational cost during on-device fine-tuning, while maintaining accuracy?
  • Q3
    Can LANCE leverage fixed low-rank subspaces to achieve competitive performance on continual learning benchmarks, while offering superior memory efficiency?
Experimental Setup.
(i) Datasets: For single-task fine-tuning, we use CIFAR-10/100 [Krizhevsky2009LearningImages], Oxford-IIIT Pets [petsdataset], Flowers-102 [flower102dataset], CUB-200 [cub200], and ImageNet [Krizhevsky2012ImageNetNetworks], covering a range of dataset sizes and complexities. For continual learning, we evaluate on Split CIFAR-100 (20 tasks), Split MiniImageNet (20 tasks), and the 5-Datasets benchmark (CIFAR-10, MNIST, SVHN, Fashion-MNIST, and notMNIST), following common CL protocols. (ii) Metrics: We report the following metrics: classification accuracy in the target datasets; memory usage, measured as peak activation storage in MB (estimated in Appendix C); and computational cost in FLOPs. For continual learning benchmarks, we additionally report the average final accuracy over all tasks (ACC) and the Backward Transfer (BWT), which quantifies the forgetting of previously learned tasks when new tasks are introduced, as defined in Appendix A.2.
Figure 2: Gradient alignment between full BP and LANCE. We plot the angle between true gradients and LANCE-projected gradients across epochs for different fine-tuning tasks. LANCE consistently produces gradients within 70 of the true gradient, and angles stabilize as training progresses, indicating preserved descent directions.
3.1 Results: Single-task Fine-Tuning Efficiency
We first evaluate LANCE on fine-tuning the last 2 or 4 layers of MCUNet [Lin2020MCUNet:Devices], MobileNetV2 [mobilenetv2], ResNet18 [he2016deep] and ResNet34 [he2016deep]. All models are pretrained on ImageNet-1k. For LANCE, the offline Phase I calibration uses N=100N{=}100 mini-batches with an energy threshold $\varepsilon{=}0.7$. Fine-tuning runs for 50 epochs.
Table 1: ImageNet-scale fine-tuning on Split ImageNet-1k. We pretrain each model on 500 ImageNet classes (Split A) and fine-tune on the remaining 500 classes (Split B).
ResNet18 ResNet34
Method Acc $\uparrow$ MB $\downarrow$ TFLOPs $\downarrow$ Acc $\uparrow$ MB $\downarrow$ TFLOPs $\downarrow$
BP (2 layers) 60.70 24.50 0.059 60.77 24.50 0.059
BP (4 layers) 64.28 61.25 0.090 64.21 49.00 0.118
LANCE (2 layers) 59.18 7.12 0.048 58.59 4.49 0.046
LANCE (4 layers) 62.20 16.59 0.075 62.66 10.79 0.093
[…]
Answer to Q2. Across all datasets and models, LANCE reduces activation storage by up to $250\times$ relative to BP while maintaining competitive accuracy. Table˜2 reports results on CIFAR-10/100 and Pets and CUB200. Compared to iterative low-rank methods such as ASI [nguyen2025beyond], or HOSVD [nguyen2024activation] where a full SVD is recomputed at each step, LANCE achieves comparable compression with higher accuracy and similar FLOPs, confirming that repeated decompositions are unnecessary. Although BP attains higher accuracy when the same number of layers are fine-tuned, LANCE’s extreme memory efficiency enables fine-tuning more layers under strict SRAM budgets. For example, fine-tuning 4 layers with LANCE reaches the accuracy of 2-layer BP while still fitting within microcontroller constraints. These results demonstrate that fixed low-rank subspaces retain sufficient representational power for efficient on-device fine-tuning.
Table 2: Performance and efficiency comparison. Results are reported in terms of accuracy, memory usage, and TFLOPs. denote the results from the respective original paper.
Mdl. Method # Layers CIFAR10 CIFAR100 Pets CUB200
Acc $\uparrow$ MB $\downarrow$ TFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ TFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ TFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ TFLOPS $\downarrow$
MCUNet BP 2 75.85 11.71 0.000 52.47 11.71 0.000 56.30 11.71 0.000 33.05 11.71 0.000
4 82.91 17.57 0.001 58.38 17.57 0.001 60.01 17.57 0.001 34.13 17.57 0.001
HOSVD 2 70.91 0.04 1.930 48.00 0.05 1.930 56.11 0.06 1.93 31.22 0.06 1.93
4 75.63 0.12 2.590 52.34 0.10 2.590 57.89 0.17 2.59 31.72 0.14 2.59
LANCE 2 70.64 0.04 0.000 47.57 0.04 0.000 57.26 0.06 0.000 31.44 0.05 0.000
4 76.23 0.09 0.000 52.59 0.09 0.000 58.05 0.14 0.000 30.87 0.13 0.000
MobileNetV2 BP 2 91.72 30.62 0.01 73.43 30.62 0.01 89.94 30.62 0.01 64.37 30.62 0.01
4 92.23 57.42 0.01 74.69 57.42 0.01 90.32 57.42 0.01 65.86 57.42 0.01
HOSVD 2 85.67 0.13 11.87 67.28 0.14 11.87 89.58 0.13 11.87 58.11 0.13 11.87
4 89.03 0.16 22.86 69.33 0.15 22.86 89.91 0.23 22.86 60.30 0.16 22.86
ASI 2 85.09 0.26 0.01 61.03 0.15 0.01 88.13 0.15 0.01 46.44 0.26 0.01
4 88.03 0.63 0.06 66.18 0.63 0.06 88.91 0.76 0.06 50.69 0.63 0.06
LANCE 2 84.16 0.15 0.01 60.69 0.16 0.01 89.69 0.14 0.01 58.18 0.14 0.01
4 88.99 0.17 0.01 68.79 0.17 0.01 90.40 0.19 0.01 61.21 0.15 0.01
ResNet18 BP 2 92.69 24.5 0.05 75.69 24.5 0.06 89.17 24.5 0.06 62.28 24.5 0.06
4 94.01 61.25 0.09 77.07 61.25 0.09 89.07 61.25 0.09 60.80 61.25 0.09
HOSVD 2 92.17 0.71 6.12 74.57 0.72 6.12 89.47 0.92 6.12 60.02 0.73 6.12
4 93.16 1.16 15.56 75.35 1.21 15.56 88.71 1.77 15.56 59.26 1.41 15.56
ASI 2 90.77 1.24 0.04 69.70 0.90 0.04 88.75 2.01 0.04 56.77 1.51 0.04
4 91.90 1.89 0.06 71.93 1.61 0.06 88.44 4.34 0.07 55.73 3.50 0.07
LANCE 2 91.8 0.75 0.03 74.09 0.77 0.03 89.26 0.94 0.03 59.87 0.73 0.03
4 92.89 1.11 0.05 75.7 01.16 0.05 89.12 1.78 0.05 59.50 1.36 0.05
ResNet34 BP 2 93.48 24.50 0.06 76.19 24.50 0.06 91.22 24.50 0.06 64.10 24.50 0.06
4 94.37 49.00 0.12 77.50 49.00 0.12 91.60 49.00 0.12 64.22 49.00 0.12
HOSVD 2 92.70 0.42 6.12 75.31 0.45 6.12 90.89 0.48 6.12 60.82 0.38 6.12
4 93.80 0.92 12.25 76.01 0.95 12.25 91.11 1.10 12.25 62.28 0.84 12.25
ASI 2 90.09 0.49 0.03 69.66 0.44 0.03 91.09 0.81 0.04 58.77 0.60 0.03
4 91.54 1.24 0.07 70.60 1.00 0.07 91.41 2.05 0.07 58.85 1.58 0.07
LANCE 2 92.62 0.46 0.03 74.89 0.51 0.03 90.70 0.45 0.03 62.35 0.39 0.03
4 93.56 1.01 0.06 75.94 1.02 0.06 90.92 1.05 0.06 63.01 0.80 0.06
Scalability to ImageNet-Scale Fine-Tuning.
[…]
Table 3: Performance comparison on continual image classification datasets using multi-head networks. Accuracy and BWT (mean $\pm$ std) are reported over five trials. denotes the results taken from [Saha2021GradientLearning] and denote the results from the respective original papers.
Method Split CIFAR100 Split MiniImageNet 5-Datasets
ACC (%) $\uparrow$ BWT (%) $\uparrow$ Mem (MB) $\downarrow$ ACC (%) $\uparrow$ BWT (%) $\uparrow$ Mem (MB) $\downarrow$ ACC (%) $\uparrow$ BWT (%) $\uparrow$ Mem (MB) $\downarrow$
Multitask $79.58\pm 0.54$ - - $69.46\pm 0.62$ - - $91.54\pm 0.28$ - -
EWC $68.80\pm 0.88$ -$2\pm 1$ - $52.01\pm 2.53$ -$12\pm 3$ - $88.64\pm 0.26$ -$4\pm 1$ -
HAT $72.06\pm 0.50$ $0\pm 0$ - $59.78\pm 0.57$ -$3\pm 0$ - $91.32\pm 0.18$ -$1\pm 0$ -
A-GEM $63.98\pm 1.22$ -$15\pm 2$ - $57.24\pm 0.72$ -$12\pm 1$ - $84.04\pm 0.33$ -$12\pm 1$ -
ER_Res $71.73\pm 0.63$ -$6\pm 1$ - $58.94\pm 0.85$ -$7\pm 1$ - $80.31\pm 0.22$ -$4\pm 0$ -
TRGP+SD $75.5\pm 0.35$ -$0.96\pm 0.09$ - $65.8\pm 0.16$ -$0.49\pm 0.08$ - - - -
LMSP $74.21$ +$0.94$ - $64.2$ +$1.55$ - $93.78$ +$0.07$ -
GPM $72.48\pm 0.54$ -$0.9$ $22.27$ $60.41\pm 0.61$ -$0.9$ $127.37$ $91.22\pm 0.20$ -$1.0$ $70.33$
TRGP $74.64\pm 0.32$ -$0.9\pm 0.01$ $78.86$ $61.78\pm 0.60$ -$0.5\pm 0.60$ $543.22$ $93.56\pm 0.10$ -$0.04\pm 0.01$ $181.57$
CUBER $75.54\pm 0.22$ $0.13\pm 0.08$ $326.79$ $62.67\pm 0.74$ $0.23\pm 0.15$ $604.61$ $93.48\pm 0.10$ -$0.00\pm 0.02$ $196.91$
SGP $76.05\pm 0.43$ -11 $22.27$ $62.83\pm 0.33$ -11 $127.37$ - - -
CODE-CL $77.21\pm 0.32$ -$1.1\pm 0.28$ $33.80$ $71.16\pm 0.32$ -$1.1\pm 0.3$ $283.82$ $93.51\pm 0.13$ -$0.11\pm 0.01$ $116.72$
LANCE $71.52\pm 0.27$ -$0.17\pm 0.34$ $2.32$ $59.68\pm 1.17$ -$0.99\pm 0.78$ $32.96$ $90.76\pm 0.31$ -$1.02\pm 0.15$ $34.96$
3.2 Results: Continual Learning
We compare LANCE against a broad set of continual learning baselines: regularization based methods (EWC [Kirkpatrick2017OvercomingNetworks], HAT [Serra2018OvercomingTask]), replay-based approaches (A-GEM [Chaudhry2019EfficientA-GEM], ER_Res [Chaudhry2019OnLearning]), and projection-based methods (GPM [Saha2021GradientLearning], TRGP [Lin2022TRGP:Learning], CUBER [Lin22CUBER], SGP [Saha2023ContinualProjection], SD [SD2023], LMSP [lmsp2025], CODE-CL [apolinario2024code]). For reference, we also report multitask performance, which serves as an upper bound where all tasks are trained jointly. We follow the same experimental setup as [apolinario2024code]: a 5-layer AlexNet for Split CIFAR-100, and a reduced ResNet18 for Split MiniImageNet and the 5-Datasets benchmark. During the first task, models are trained without constraints using full BP. For subsequent tasks, the model from the previous task serves as initialization, and LANCE applies its subspace-constrained training.
Answer to Q3. LANCE achieves accuracy competitive with full-rank gradient projection methods such as GPM and CODE-CL, while using drastically less memory. As summarized in Tab.˜3, on Split CIFAR-100 LANCE reaches $\sim$71% average accuracy (ACC) while reducing activation storage by more than $10\times$ compared to GPM. Similar trends are observed on Split MiniImageNet and the 5-Datasets benchmark, where LANCE maintains stable performance without rehearsal memory. These results confirm that fixed low-rank subspaces are sufficient to retain representational capacity across tasks, while providing substantial memory savings for resource-constrained devices.
Figure 3: Ablation studies of LANCE. (Left) Effect of the energy threshold $\varepsilon$ on CUB-200 using MCUNet and ResNet34. Accuracy improves steadily as $\varepsilon$ increases, but memory usage grows exponentially, illustrating the trade-off between accuracy and compression. (Right) Effect of the number of calibration batches NN on CIFAR-100. Accuracy remains stable even for very small NN, while memory increases with larger calibration sets. These results show that stable subspaces can be obtained with as few as N=2N{=}2 calibration batches, and practical choices of $\varepsilon$ (e.g., 0.7) provide a good balance between accuracy and memory.
3.3 Ablations and Analysis
We conduct ablations on two key hyperparameters: (i) the energy threshold $\varepsilon$ (evaluated on CUB-200), and (ii) the number of calibration batches $N$ (evaluated on CIFAR-100 with $\varepsilon{=}0.7$). Results for MCUNet and ResNet34 are shown in Figure˜3. As expected, higher $\varepsilon$ increases accuracy but also grows memory usage exponentially, highlighting the need to balance compression and accuracy according to device memory budgets. For calibration size $N$, accuracy remains stable across a wide range, while memory usage grows with $N$. Remarkably, even N=2N{=}2 calibration batches are sufficient to yield stable subspaces, underscoring the robustness of the one-shot calibration phase.
Figure 4: Gradient fidelity across compression ratios. We vary the energy threshold $\varepsilon$ controlling the effective memory compression of LANCE and report the angle (mean $\pm$ std) between full-BP gradients and LANCE-projected gradients over the final 10 epochs (out of 50). As $\varepsilon$ decreases, compression becomes more aggressive and gradient alignment degrades accordingly. Nevertheless, even under memory reductions of up to two orders of magnitude, LANCE maintains gradient directions within $\sim 70^{\circ}$ of full BP, indicating preserved descent directions despite strong compression.
Figure 5: End-to-end fine-tuning latency on a Raspberry Pi 3B+ (MCUNet on CIFAR-10, batch size 128). Averaged over five trials, LANCE consistently achieves the lowest forward, backward, and total training times across all layer depths.
We additionally evaluate how different compression ratios affect gradient fidelity by measuring the angle between full-BP gradients and LANCE-projected gradients across several values of $\varepsilon$, shown in Figure˜4. As expected, more aggressive compression (lower $\varepsilon$) leads to larger angles, but the degradation is smooth and controlled. Even at the strongest compression settings, reducing activation storage by up to two orders of magnitude, LANCE maintains gradient alignment within $\sim 70^{\circ}$ of full BP, indicating that the one-shot subspace preserves the dominant descent directions required for stable optimization.
3.4 Runtime Validation on Edge Hardware
[…]
4 Related Work
Storing intermediate activations is the primary memory bottleneck for on-device training [lin2020tinytl, lin2022ondevice]. Prior work has reduced this cost through recomputation (e.g., checkpointing [chen2016checkpoint]), architectural modifications (e.g., reversible networks [gomez2017reversible]), or lossy compression and sparsity [chen2021actnn, barley2023activation, yu2022backrazor], but often at the expense of extra computation, accuracy, or hardware complexity. A complementary line of work exploits low-rank structure in activations [nguyen2024activation, nguyen2025beyond]. Closest to our method is ASI [nguyen2025beyond], which incrementally updates subspaces during training; in contrast, LANCE performs a one-shot HOSVD at initialization and reuses the resulting subspaces throughout fine-tuning, avoiding repeated decompositions. Recent CL work such as LMSP [lmsp2025] also uses low-rank representations but focuses on reducing the computational cost of gradient projection, still requiring multiple per-task low-rank memory components; by contrast, LANCE targets the dominant activation-memory bottleneck directly. This simple design not only reduces compute and hardware overhead but also enables a continual learning extension: fixed subspaces naturally support null-space allocation across tasks without storing large task-specific matrices. System-level methods such as TinyTL [lin2020tinytl], MCUNet [Lin2020MCUNet:Devices], and the 256KB training engine [lin2022ondevice] co-design architectures and update rules to fit extreme SRAM budgets, often by freezing layers or pruning gradient paths. LANCE is complementary to these approaches: instead of restricting the optimization, it preserves full backpropagation while directly compressing activations. Finally, in continual learning, regularization methods [Kirkpatrick2017OvercomingNetworks, Serra2018OvercomingTask], replay-based approaches [Chaudhry2019EfficientA-GEM, Chaudhry2019OnLearning], and gradient-projection methods [Saha2021GradientLearning, Lin2022TRGP:Learning, Lin22CUBER, Saha2023ContinualProjection, apolinario2024code, SD2023] all mitigate forgetting by controlling parameter updates or storing exemplars. LANCE differs by operating directly in the activation subspace: new tasks are assigned to the orthogonal complement of previously used directions during calibration. This yields performance competitive with gradient-projection methods while requiring significantly less memory, and without rehearsal buffers.
5 Conclusion
We introduced LANCE, a one-shot low-rank activation compression framework for efficient on-device learning. By calibrating reusable activation subspaces with a single HOSVD, LANCE eliminates repeated decompositions and reduces both memory and computational cost. Experiments show that LANCE achieves up to $250\times$ memory savings while maintaining accuracy close to full backpropagation across diverse datasets and models. Crucially, the fixed low-rank subspaces naturally extend to continual learning, where LANCE matches the performance of gradient projection methods at a fraction of their memory footprint. Together, these results establish LANCE as a practical and scalable approach for enabling fine-tuning and continual learning on resource-constrained edge devices.
Acknowledgments
This work was supported in part by the Center for Co-design of Cognitive Systems (CoCoSys), one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program, and in part by the Department of Energy (DoE).
Appendix 0.A Datasets, Models and Hyperparameters
This section provides details on the architecture of all models used in this work, the dataset statistics, the hyperparameters for each experiment, and the compute resources employed.
Reported metrics.
For single-task fine-tuning we report top-1 accuracy on the test split, peak activation memory in MB (estimated from tensors retained for the backward pass), and training FLOPs. [^1] Peak activation memory and FLOPs formulas are detailed in Appendix 0.C. For continual learning experiments (Section 2.3), we additionally report ACC and BWT following standard CL definitions.
0.A.1 For Single-Task Fine-Tuning Experiments
0.A.1.1 Datasets
We evaluate on five image classification datasets spanning scale and difficulty: CIFAR-10 [Krizhevsky2009LearningImages], CIFAR-100 [Krizhevsky2009LearningImages], Oxford-IIIT Pets [petsdataset], Flowers-102 [flower102dataset], and CUB-200 [cub200]. We use the official training/test splits provided by each dataset and do not carve out an additional validation set. For all datasets, images are resized to match the native input resolution of the chosen architecture (e.g., $224\times 224$ for ImageNet-pretrained models), and standard data augmentation is applied consisting of random horizontal flipping.
0.A.1.2 Model Architectures
For single-task experiments we use four pretrained models: ResNet18, ResNet34 [he2016deep], MobileNetV2 [mobilenetv2], and MCUNet [Lin2020MCUNet:Devices]. All of them are initialized from ImageNet-1k pretrained weights. For each model, we unfreeze the last $N\in\{2,4\}$ convolutional layers and fine-tune them while the rest of the network remains frozen.
0.A.1.3 Training Protocol
Training runs for 50 epochs with stochastic gradient descent (SGD), learning rate 0.05, batch size 128, and without weight decay. For LANCE, we allocate N=100N{=}100 calibration mini-batches to estimate the activation covariance (unless otherwise specified) and compute truncated subspaces. The energy threshold $\varepsilon$ for subspace retention $0.7$. During training, activations of the selected layers are projected into the fixed low-rank cores, which are the only tensors retained for the backward pass.
0.A.1.4 Hyperparameters
The hyperparameters used in our single-task fine-tuning experiments are summarized in Table 4.
Table 4: Hyperparameters used in single-task fine-tuning experiments.
Hyperparameter Value
Epochs 50
Optimizer SGD
Learning rate 0.05
Batch size 128
Weight decay 0
# Unfrozen layers 2 or 4
# Calibration batches ($N$) 100
LANCE threshold $\varepsilon$ 0.7
0.A.2 For Continual Learning Experiments
Table 5: 5-Datasets statistics.
Dataset CIFAR10 MNIST SVHN Fashion MNIST notMNIST
Number of classes 10 10 10 10 10
Training samples 47500 57000 69595 57000 16011
Validation samples 2500 3000 3662 3000 842
Test samples 10000 10000 26032 10000 1873
Table 6: Split CIFAR100 and Split miniImageNet datasets statistics.
Dataset Split CIFAR100 Split miniImageNet
Number of tasks ($T$) 10 20
Sample dimensions $3\times 32\times 32$ $3\times 84\times 84$
Number of classes per task 10 5
Training samples per task 4750 2375
Validation samples per task 250 125
Test samples per task 1000 500
Table 7: List of hyperparameters used in Continual Learning experiments.
Dataset Split CIFAR100 Split miniImageNet 5-Datasets
Optimizer SGD SGD SGD
Learning rate ($\eta$) $0.01$ $0.1$ $0.1$
Batch size ($b$) $64$ $64$ $64$
Min. learning rate ($\eta_{th}$) $10^{-5}$ $10^{-5}$ $10^{-3}$
Learning rate decay factor $1/2$ $1/2$ $1/3$
Patience 66 66 55
Number of epochs ($E$) $200$ $100$ $100$
Energy Threshold ($\varepsilon$) $0.9$ $0.9$ $0.9$
CL Threshold ($\varepsilon_{\text{CL}}$) $0.95$ $0.985$ $0.97$
# Calibration batches ($N$) $100$ $100$ $100$
# batches for memory retention ($N_{\text{CL}}$) $10$ $10$ $10$
0.A.2.1 Model Architecture
We employ two neural network architectures in our experiments: an AlexNet-inspired model following [Serra2018OvercomingTask], and a Reduced ResNet18 as described in [Lopez-Paz2017GradientLearning]. The AlexNet-like network integrates batch normalization (BN) layers after every convolutional and fully connected layer except the final classifier. BN parameters are updated only during the first task and kept fixed thereafter. Its architecture consists of three convolutional layers with $64$, $128$, and $256$ filters, using kernel sizes of $4\times 4$, $3\times 3$, and $2\times 2$, respectively, each followed by a $2\times 2$ max-pooling operation. Two fully connected layers with $2048$ units each are appended. ReLU activations are applied throughout, and dropout is used with rates of $0.2$ for the first two layers and $0.5$ for the rest. The Reduced ResNet18 follows the setup in [Saha2021GradientLearning]. For Split miniImageNet tasks, its first convolutional layer uses stride $2$, whereas for the 5-Datasets benchmark, stride $1$ is used. Across all experiments, models are trained with cross-entropy loss.
0.A.2.2 Dataset Statistics
Tables 6 and 5 summarize the datasets employed in our continual learning benchmarks. We adopt the same train/test splits as in [Saha2021GradientLearning, Lin2022TRGP:Learning, Saha2023ContinualProjection, apolinario2024code]. For the 5-Datasets benchmark, grayscale samples are duplicated across three channels to match RGB input requirements. All images are resized to $32\times 32$, yielding an effective input dimension of $3\times 32\times 32$.
0.A.2.3 Metrics
Following standard practice in prior continual learning works [Lopez-Paz2017GradientLearning, Saha2021GradientLearning, Lin2022TRGP:Learning, Saha2023ContinualProjection, apolinario2024code], we evaluate performance using two metrics: the average final accuracy across all tasks (ACC) and the Backward Transfer (BWT), which captures the extent to which learning new tasks affects performance on previously learned ones. They are defined as:
$\text{ACC}=\sum_{i=1}^{T}\frac{{A}_{T,i}}{T}\text{ ; }\text{BWT}=\sum_{i=1}^{T-1}\frac{{A}_{T,i}-{A}_{i,i}}{T-1},$ (6)
where $T$ denotes the total number of tasks, and ${A}_{j,i}$ represents the accuracy on the $i$-th task after the model has sequentially trained up to the $j$-th task $(i\leq j)$.
0.A.2.4 Hyperparameters
The hyperparameters used in our experiments are detailed in Table 7.
0.A.3 Compute resources
All experiments were conducted on a shared internal Linux server equipped with an AMD EPYC 7502 32-Core Processor, 504 GB of RAM, and four NVIDIA A40 GPUs, each with 48 GB of GDDR6 memory. Additionally, code was implemented using Python 3.9 and PyTorch 2.2.1 with CUDA 11.8.
Appendix 0.B Proofs for Section 2.4
See 2.1
Proof
Let $\delta$ denote the upstream error for the layer and let $\bm{x}$} be the unfolded input. For a linear (or unfolded convolutional) layer, the full gradient is $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}=\delta\,{\bm{x}}^{\top}$. In LANCE, the only change (by design) is replacing $\bm{x}$} with $\mathsf{P}^{(l)}{\bm{x}}$, so
$\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}=\delta\,(\mathsf{P}^{(l)}{\bm{x}})^{\top}=\delta\,{\bm{x}}^{\top}\mathsf{P}^{(l)}=\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\,\mathsf{P}^{(l)}.$
Since $\mathsf{P}^{(l)}$ is an orthogonal projector (Assumption A1),
$\big\langle\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}},\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\big\rangle$ $=\mathrm{tr}\!\left(\big(\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\mathsf{P}^{(l)}\big)^{\top}\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\right)$
$=\|\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{full}}\mathsf{P}^{(l)}\|_{F}^{2}$
$=\|\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}\|_{F}^{2}\geq 0.$
Thus $\nabla_{{\bm{W}}}\mathcal{L}_{\mathrm{LANCE}}$ is a descent direction unless it is zero.
See 2.2
Proof
By the Descent Lemma for an $L$-smooth function,
$\mathcal{L}({\bm{\theta}}-\eta{\bm{g}})\leq\mathcal{L}({\bm{\theta}})-\eta\langle\nabla\mathcal{L}({\bm{\theta}}),{\bm{g}}\rangle+\tfrac{L\eta^{2}}{2}\|{\bm{g}}\|_{2}^{2}.$
Set ${\bm{g}}=\mathsf{Q}\,\nabla\mathcal{L}({\bm{\theta}})$. Because $\mathsf{Q}$ is an orthogonal projector (blockwise right-multiplication by $\mathsf{P}^{(l)}$), $\langle\nabla\mathcal{L},\mathsf{Q}\nabla\mathcal{L}\rangle=\|\mathsf{Q}\nabla\mathcal{L}\|_{2}^{2}$. Hence
$\mathcal{L}({\bm{\theta}}_{k+1})\leq\mathcal{L}({\bm{\theta}}_{k})-\eta\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}+\tfrac{L\eta^{2}}{2}\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}\leq\mathcal{L}({\bm{\theta}}_{k})-\tfrac{\eta}{2}\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}$
for $\eta\leq 1/L$. Summation gives $\sum_{k}\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}^{2}<\infty$ and thus $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Any limit point $\bm{\theta}$$_{\star}$ satisfies $\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{\star})=\mathbf{0}$.
See 1
Proof
From Theorem˜2.2, $\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\to 0$. Then
$\|\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\;\leq\;\|\mathsf{Q}\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}+\|(I-\mathsf{Q})\nabla\mathcal{L}({\bm{\theta}}_{k})\|_{2}\;\to\;\leq\;C\sqrt{1-\varepsilon}.$
Take a limit point $\bm{\theta}$$_{\star}$.
Remark 1(Why the leakage bound is reasonable)
Right-multiplication by $\mathsf{P}^{(l)}$ removes only the gradient components aligned with the discarded input subspaces. One-shot HOSVD calibration bounds the relative energy of those components by $\sqrt{1-\varepsilon}$ in Frobenius norm, and layer-wise Lipschitz constants contract this via the chain rule to produce the stated bound.
Appendix 0.C Complexity Analysis of LANCE
Let ${\bm{\mathsfit{X}}}^{(l)}\in{\mathbb{R}}^{n_{1}\times n_{2}\times n_{3}}$, its compressed core ${\bm{\mathsfit{G}}}^{(l)}\in{\mathbb{R}}^{r_{1}\times r_{2}\times r_{3}}$, and the next-layer output ${\bm{\mathsfit{X}}}^{(l+1)}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\hat{n}_{3}}$. Assume parameters ${\bm{\theta}}^{(l)}\in{\mathbb{R}}^{n_{3}\times\hat{n}_{3}}$ operate on the last dimension. We estimate FLOPs and memory during fine-tuning and omit the one-shot HOSVD cost since it is performed offline.
The forward compression cost is $\sum_{i=1}^{3}n_{i}r_{i}\prod_{j\neq i}r_{j}$, the cost of projecting ${\bm{\mathsfit{X}}}^{(l)}$ with $\{{\bm{U}}^{(l)}_{i}\}_{i=1}^{3}$. Adding the layer’s standard forward cost $\hat{n}_{3}\prod_{i=1}^{3}n_{i}$ gives:
$\mathrm{FLOPs}^{\mathrm{FW}}_{\mathrm{LANCE}}\;=\;\hat{n}_{3}\prod_{i=1}^{3}n_{i}\;+\;n_{1}r_{1}n_{2}n_{3}\;+\;r_{1}n_{2}r_{2}n_{3}\;+\;r_{1}r_{2}n_{3}r_{3}.$ (7)
Similarly, the backward pass FLOPs with LANCE can be estimated as:
$\mathrm{FLOPs}^{\mathrm{BW}}_{\mathrm{LANCE}}\;=\;\prod_{i=1}^{3}n_{i}r_{1}\;+\;\prod_{i=1}^{3}r_{i}n_{2}\;+\;r_{1}n_{2}r_{3}n_{3}\;+\;r_{1}n_{2}n_{3}\hat{n}_{3}.$ (8)
For full BP, the forward and backward FLOPs are:
$\mathrm{FLOPs}^{\mathrm{FW}}_{\mathrm{BP}}\;=\;\hat{n}_{3}\prod_{i=1}^{3}n_{i},\quad\mathrm{FLOPs}^{\mathrm{BW}}_{\mathrm{BP}}\;=\;\hat{n}_{3}\prod_{i=1}^{3}n_{i}.$ (9)
Therefore, the overall improvement is
$S_{\mathrm{FLOPs}}\;=\;\frac{\mathrm{FLOPs}^{\mathrm{FW}}_{\mathrm{BP}}+\mathrm{FLOPs}^{\mathrm{BW}}_{\mathrm{BP}}}{\mathrm{FLOPs}^{\mathrm{FW}}_{\mathrm{LANCE}}+\mathrm{FLOPs}^{\mathrm{BW}}_{\mathrm{LANCE}}}.$
Since $r_{i}\ll n_{i}$, we obtain an improvement over BP on the order of $\sim 1.5\times$, consistent with Fig. 1c.
For memory usage, the improvement is measured as
$S_{\mathrm{Mem}}\;=\;\frac{\mathrm{Mem}_{\mathrm{BP}}}{\mathrm{Mem}_{\mathrm{LANCE}}}\;=\;\frac{\prod_{i=1}^{3}n_{i}}{\prod_{i=1}^{3}r_{i}+\sum_{i=1}^{3}n_{i}r_{i}}.$
Thus, LANCE reduces activation storage by a factor of $S_{\mathrm{Mem}}$ with negligible amortized overhead, since the decomposition is one-shot rather than repeated every iteration.
Appendix 0.D Additional Gradient-Compression Ablations
For completeness, we extend the gradient fidelity analysis of Sec. 3.3 to three additional fine-tuning datasets: Flowers102, Oxford-IIIT Pets, and CIFAR-10. Using the same experimental setup as in the main paper, we vary the energy threshold $\varepsilon$ to induce different activation compression ratios and measure (i) the angle between full-BP gradients and LANCE-projected gradients, (ii) the resulting memory ratio, and (iii) the accuracy obtained under each setting. Appendix Fig. 6 shows results for MCUNet and ResNet34. The behavior is highly consistent with the CUB-200 and CIFAR-100 experiments: gradient alignment degrades smoothly as compression increases, yet even aggressive compression ($\sim$2 orders of magnitude memory reduction) preserves descent directions within $70^{\circ}$. These findings demonstrate that the one-shot activation subspace generalizes well across models and datasets, further validating the stability and robustness of LANCE’s low-rank representations.
Figure 6: Additional gradient–compression ablations on Flowers, Pets, and CIFAR-10. We repeat the analysis of Fig. 4 on three additional fine-tuning datasets using MCUNet and ResNet34. For each dataset, we report the mean and standard deviation of the angle between full-BP gradients and LANCE-projected gradients over the final 10 epochs of training, together with the corresponding memory ratio (LANCE/BP) and accuracy across different energy thresholds $\varepsilon$. Across all datasets and architectures, gradient alignment degrades smoothly as $\varepsilon$ decreases, while extreme compression (up to two orders of magnitude in activation memory) continues to yield descent directions within $\sim 70^{\circ}$ of full BP. These trends closely mirror those observed in the main experiments, reinforcing the robustness and consistency of LANCE’s low-rank subspace across diverse data distributions.
Appendix 0.E Robustness of LANCE Under Small-Batch Regimes
To evaluate LANCE under more restrictive memory settings, we additionally benchmarked batch sizes $\text{BS}=4$ and $\text{BS}=8$, which are more representative of MCU- and IoT-class devices. The results (Tables 89) show that LANCE maintains the same qualitative behavior observed with $\text{BS}=128$: memory reductions of 1–2 orders of magnitude, negligible FLOPs overhead, and accuracy comparable to full BP. This indicates that LANCE’s efficiency gains do not rely on large batch sizes and extend reliably to resource-constrained edge settings.
Table 8: Performance and efficiency comparison. Results are reported in terms of accuracy, memory usage, and Giga FLOPS (GFLOPs) for different batch sizes (BS).
Mdl. Method # Layers CIFAR10 (BS = 4) CIFAR10 (BS = 8) CIFAR100 (BS = 4) CIFAR100 (BS = 8)
Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$
MCUNet BP 2 77.17 0.220 0.025 76.87 0.439 0.049 54.24 0.220 0.025 53.96 0.439 0.049
4 83.93 0.403 0.039 83.68 0.806 0.079 60.99 0.403 0.039 59.27 0.806 0.079
LANCE 2 69.72 0.007 0.014 76.32 0.011 0.028 45.70 0.008 0.014 51.27 0.010 0.028
4 74.23 0.018 0.023 77.89 0.020 0.045 49.23 0.015 0.023 53.36 0.019 0.045
ResNet34 BP 2 94.08 0.766 1.850 93.92 1.531 3.699 77.85 0.766 1.850 77.99 1.531 3.699
4 94.96 1.531 3.699 95.10 3.062 7.399 79.02 1.531 3.699 79.10 3.062 7.399
LANCE 2 91.75 0.059 1.018 92.29 0.093 2.022 73.36 0.061 1.021 74.51 0.101 2.036
4 92.05 0.168 2.114 93.20 0.232 4.124 73.09 0.156 2.096 76.17 0.252 4.161
Table 9: Performance and efficiency comparison. Results are reported in terms of accuracy, memory usage, and Giga FLOPS (GFLOPs) for different batch sizes (BS).
Mdl. Method # Layers CUB200 (BS = 4) CUB200 (BS = 8) PETS (BS = 4) PETS (BS = 8)
Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$ Acc $\uparrow$ MB $\downarrow$ GFLOPS $\downarrow$
MCUNet BP 2 32.57 0.220 0.025 32.84 0.439 0.049 55.52 0.220 0.025 55.55 0.439 0.049
4 37.50 0.403 0.039 37.04 0.806 0.079 61.98 0.403 0.039 61.35 0.806 0.079
LANCE 2 34.79 0.011 0.015 34.97 0.013 0.029 60.32 0.011 0.015 59.85 0.015 0.029
4 35.90 0.024 0.024 36.12 0.029 0.047 60.75 0.028 0.024 59.72 0.032 0.047
ResNet34 BP 2 66.50 0.766 1.850 66.86 1.531 3.699 91.44 0.766 1.850 91.11 1.531 3.699
4 69.09 1.531 3.699 69.57 3.062 7.399 91.14 1.531 3.699 91.63 3.062 7.399
LANCE 2 65.53 0.053 1.006 64.81 0.094 2.022 91.20 0.075 1.043 90.87 0.111 2.055
4 65.72 0.160 2.102 66.31 0.220 4.102 90.84 0.182 2.136 90.92 0.275 4.202
Appendix 0.F Limitations and Future Direction.
LANCE opens several promising directions for future work. First, while the method relies on activation subspaces obtained from a small calibration set, developing mechanisms to adapt or refine these subspaces under strong domain shift or evolving representations would broaden its applicability—including extending the approach to earlier stages of pre-training, where representation drift is more pronounced. In continual learning, relaxing the strict orthogonality constraint could enable positive forward transfer between related tasks while retaining stability. Additionally, augmenting the fixed low-rank subspace with selective or lightweight updates may improve adaptability on tasks requiring larger representational changes. Finally, although LANCE is compatible with MCU-class devices, integrating it with device-specific compilation and memory-scheduling pipelines presents an exciting opportunity to realize its full hardware potential.
Appendix 0.G Code Availability
The implementation of LANCE, together with scripts to reproduce all experiments, will be released in a public repository after the review process. This will include training and evaluation code, dataset preprocessing pipelines, and experiment configuration files to ensure full reproducibility.